In a public telephone booth the arrivals are on the average 15 per hour. A call on the average takes 3minutes.If there is just one phone,find

a)expected number of callers in the booth at anytime

b)The portion of the time the booth is expected to be idle


1
Expert's answer
2021-03-25T06:33:15-0400

"\lambda=\frac{1}{15},"

"\mu=\frac 13,"

1)

"L=\frac{\mu}{\mu-\lambda}=\frac{\frac 13}{\frac 13-\frac{1}{15}}=1.25,"

2)

"P=1-\frac{\lambda}{\mu}=1-\frac{\frac{1}{15}}{\frac 13}=0.8."